OpenStudy (anonymous):

Dy/dx=6+sqrt{y-6x+5}

OpenStudy (anonymous):

In the future, please state exactly what you want to solve instead of only the problem itself. Assuming you want to find the derivative and assuming I can treat the problem like it's a partial derivative (aka I can treat y as a constant) : let u = y - 6x + 5 du = -6 dy/dx = 6 - (1/6)(u)^1/2 du Integrate both sides y = 6x -(1/9)(u)^(3/2) y = 6x -(1/9)(y - 6x + 5)^3/2

OpenStudy (anonymous):

Looks more like a differential equation to me: \[\frac{dy}{dx}=6-\sqrt{y-6x+5}\] Let's try a substitution, \(v=y-6x\), so that \(\dfrac{dv}{dx}=\dfrac{dy}{dx}-6\): \[\begin{align*} \frac{dv}{dx}+6&=6-\sqrt{v+5}\\ \frac{dv}{dx}&=-\sqrt{v+5}\\ \frac{dv}{\sqrt{v+5}}&=-dx \end{align*}\]